# Optimal Multi-Distribution Learning

### v2

@@ 2 Problem formulation
This section formulates the multi-distribution learning problem. We assume throughout that each datapoint takes the form of $(x,y)\in\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ (resp. $\mathcal{Y}$) representing the instance space (resp. label space).
@@
The problem setting encompasses several elements as follows.
  • Hypothesis class. Suppose that we are interested in a hypothesis class $\mathcal{H}$, consisting of a set of candidate functions from the instance space $\mathcal{X}$ to the label space $\mathcal{Y}$. Overloading the notation, we use $h_{\pi}$ to represent a randomized hypothesis associated with a probability distribution $\pi\in\Delta(\mathcal{H})$, meaning that a hypothesis $h$ from $\mathcal{H}$ is randomly selected according to distribution $\pi$. Additionally, the Vapnik–Chervonenkis (VC) dimension (Vapnik et al.,, 1994) of $\mathcal{H}$ is assumed to be
    $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d;$ (6)
    see Shalev-Shwartz and Ben-David, (2014) for an introduction of the VC dimension.
@@ 3 Algorithm
In this section, we present our proposed algorithm. Before proceeding, we find it convenient to introduce some notation concerning the loss under mixed distributions. Specifically, for any distribution $w=[w_{i}]_{1\leq i\leq k}\in\Delta(k)$ and any hypothesis $h\in\mathcal{H}$, the risk over the mixture $\sum_{i\in[k]}w_{i}\mathcal{D}_{i}$ of data distributions is denoted by:
@@
2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}:=\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}.$
@@ 9                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{n_{i}^{t}}\cdot\sum_{j=1}^{n_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (9)
@@
             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
@@
Algorithm 1 Hedge for Multi-distribution Learning ($\mathtt{MDL-Hedge}$)
Following the game dynamics proposed in previous works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), our algorithm alternates between computing the most favorable hypothesis (performed by the learner) and estimating the most challenging mixture of data distributions (performed by the adversary), with the aid of no-regret learning algorithms (Roughgarden,, 2016; Shalev-Shwartz,, 2012). More specifically, in each round $t$, our algorithm performs the following two steps:
@@
    Given hypothesis $h^{t}$, we compute an updated weight vector $w^{t+1}\in\Delta(k)$ — and hence an updated mixed distribution $\mathcal{D}^{(t+1)}=\sum_{i\in[k]}w_{i}^{t+1}\mathcal{D}_{i}$. The weight updates are carried out using the celebrated Hedge algorithm (Freund and Schapire,, 1997) designed for online adversarial learning, [^4] Note that the Hedge algorithm is closely related to Exponentiated Gradient Descent, Multiplicative Weights Update, Online Mirror Descent, etc (Arora et al.,, 2012; Shalev-Shwartz,, 2012; Hazan,, 2022). in an attempt to achieve low regret even when the loss vectors are adversarially chosen. More precisely, we run
@@
At the end of the algorithm, we output a randomized hypothesis $h^{\sf final}$ that is uniformly distributed over the hypothesis iterates $\{h^{t}\}_{1\leq t\leq T}$ over all $T$ rounds, following common practice in online adversarial learning.
While the above paradigm has been adopted in past works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), the resulting sample complexity depends heavily upon how data samples are collected and utilized throughout the learning process. For instance, Awasthi et al., (2023, Algorithm 1) draws fresh data at each step of every round, in order to ensure reliable estimation of the loss function of interest through elementary concentration inequalities. This strategy, however, becomes wasteful over time, constituting the main source of its sample sub-optimality.
@@
  • Sample reuse in Step (a). In stark contrast to Awasthi et al., (2023, Algorithm 1) that draws new samples for estimating each $h^{t}$, we propose to reuse all samples collected in Step (a) up to the $t$-th round to assist in computing $h^{t}$. As will be made precise in lines 1-1 of Algorithm 1, we shall maintain a growing dataset $\mathcal{S}$ for conducting Step (a) throughout, ensuring that there are $n_{i}^{t}$ samples drawn from distribution $\mathcal{D}_{i}$ in the $t$-th round. These datapoints are employed to construct an empirical loss estimator $\widehat{L}^{t}(h,w^{t})$ for each $h\in\mathcal{H}$ in each round $t$, with the aim of achieving uniform convergence $|\widehat{L}^{t}(h,w^{t})-L(h,w^{t})|\leq O(\varepsilon)$ over all $h\in\mathcal{H}$. More detailed explanations are provided in Section 4.1.
  • Weighted sampling for Step (b). As shown in line 1 of Algorithm 1, in each round $t$, we sample each $\mathcal{D}_{i}$ a couple of times to compute the empirical estimator for $\mathbb{E}_{(x,y)\in\mathcal{D}_{i}}\big{[}\ell(h^{t},(x,y))\big{]}$, where the number of samples depends upon the running weights $\{w_{i}^{\tau}\}$. More precisely, we collect $\left\lceil k\overline{w}_{i}^{t}\right\rceil$ fresh samples from each $\mathcal{D}_{i}$, where $\overline{w}_{i}^{t}:=\max_{1\leq\tau\leq t}w_{i}^{\tau}$ is the maximum weight assigned to $\mathcal{D}_{i}$ up to now. Informally speaking, this strategy ensures reduced variance of the estimators and ultimately allows for an improved bound for $\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}$. The interested reader is referred to Section 4.2 and Lemma 29 for more detailed explanations.
@@
4 Technique overview
In this section, we introduce the high-level ideas of the algorithm and analysis. Since Haghtalab et al., (2022) provided a nearly matching upper bound when $k=O(1/\varepsilon)$ (See the second row in Table 1), we assume that $k\geq\frac{100}{\varepsilon}$ throughout this work. We discuss the two primary technical novelties of the uniform convergence of the importance sampling estimator that allows for reusing samples (Section 4.1), and the upper bound of the coordinate-wise maximal $\ell_{1}$ norm of the $\{w^{t}\}$ iterates (Section 4.2).
4.1 Uniformly-convergent estimators
Let $\mathcal{X}$ be a proper discretization of $\Delta(k)$ with size $\exp(O(k\log(kd/\varepsilon)))$. We consider to study the approximation of the natural importance sampling estimator. Fix $n=\{n_{i}\}_{i=1}^{k}$ where each $n_{i}$ is a positive integer. Suppose that we have $n_{i}$ i.i.d. samples from $\mathcal{D}_{i}$ for each $i\in[k]$, denoted by $\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$. For $w=[w_{i}]_{1\leq i\leq k}\in\mathcal{X}$, we let
$\widehat{L}_{n}(h,w)\coloneqq\sum_{i=1}^{k}w_{i}\cdot\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}.$ (12)
In view of the concentration property of VC classes (Mohri et al.,, 2018), we have
$\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\exp\left(\lambda\cdot\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\right)\right]$
$\leq 2\cdot(2\|n\|_{1})^{d+1}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$ (13)
for all $0\leq\lambda\leq\min_{i}\frac{n_{i}}{w_{i}}$. As a result, for any $\kappa\geq 0$ and $0\leq\lambda\leq\kappa$ one has
$\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\exp\left(\lambda\cdot\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\right)\right]$
$\leq\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\cdot 2\cdot(2\|n\|_{1})^{d+1}\exp\left(2\lambda^{2}\sum_{i=1}^{k}\frac{(w_{i})^{2}}{n_{i}}\right)$
$\leq\sum_{w\in\mathcal{X}}\mathds{1}\left\{\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}\cdot 2\cdot(2\|n\|_{1})^{d+1}\exp\left(\frac{2\lambda^{2}}{\kappa}\right)$
$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\exp\left(\frac{2\lambda^{2}}{\kappa}\right).$ (14)
We can then readily see that: for any $\varepsilon^{\prime}\in(0,1)$,
@@ 4.1 Uniformly-convergent estimators
$\mathbb{P}\left\{\exists w\in\mathcal{X},\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa,\max_{h\in\mathcal{H}}\left|\widehat{L}_{n}(h,w)-L(h,w)\right|\geq\varepsilon^{\prime}\right\}$
$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\min_{0\leq\lambda\leq\kappa}\exp\left(-\lambda\varepsilon^{\prime}+\frac{2\lambda^{2}}{\kappa}\right)$
$\leq 2\cdot|\mathcal{X}|\cdot(2\|n\|_{1})^{d+1}\exp\left(-\frac{\kappa(\varepsilon^{\prime})^{2}}{4}\right).$ (15)
Choose $\varepsilon^{\prime}=c\varepsilon$ for some proper constant $c$. By virtue of (15), for any $\delta\in(0,1)$, we can choose $\kappa\geq N_{0}(\varepsilon,\delta):=100\frac{k\log(kd/\varepsilon)+d\log(\|n\|_{1})+\log(1/\delta)}{\varepsilon^{2}}$ to ensure (15) is bounded by $\delta$. In other words, with probability at least $1-\delta$, $|\widehat{L}_{n}(h,w)-L(h,w)|\leq c\varepsilon$ holds uniformly for all $w\in\mathcal{X}$ such that $\max_{i}\frac{w_{i}}{n_{i}}\leq\frac{1}{N_{0}(\varepsilon,\delta)}$.
In the analysis above, we fix $n$ and show that uniform convergence holds with high probability. In fact, with similar arguments we can obtain uniform convergence over all proper choices of $(n,w)$. Let
[…]
$\mathcal{L}(N,\kappa)$$\coloneqq\left\{(w,n)\mid w\in\mathcal{X},n\in\mathcal{N}(N),\min_{i}\frac{n_{i}}{w_{i}}\geq\kappa\right\}.$ (16b)
It can then be shown that for any $\varepsilon^{\prime}\in(0,1)$ and $\kappa>0$,
@@
$\mathbb{P}\left[\exists(w,n)\in\mathcal{L}(N,\kappa),\left|\widehat{L}_{h}(h,w)-L(h,w)\right|\geq\varepsilon^{\prime}\right]$
$\leq 2\cdot|\mathcal{N}(N)|\cdot|\mathcal{X}|\cdot(2N)^{d+1}\exp\left(-\frac{\kappa(\varepsilon^{\prime})^{2}}{4}\right).$
Since $|\mathcal{N}(N)|\leq N^{k}$, it suffices to choose $\kappa=N^{*}\coloneqq O\left(\frac{(k+d)\mathrm{polylog}(k,d,1/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}\right)$ to ensure that the $O(\varepsilon)$-uniform convergence holds with probability at least $1-\delta$.
@@
With the arguments above, with probability at least $1-\delta$, whenever we collect $n=\{n_{i}\}_{i=1}^{k}$ samples in the learning process, we could obtain $\varepsilon$-approximation $\widehat{L}_{n}(h,w)$ (see (12)) of $L(h,w)$ for all $h\in\mathcal{H}$ and $w\in\Delta(k)$ such that $\min_{i}\frac{n_{i}}{w_{i}}\geq N^{*}$. Most importantly, this result allows us to reuse samples in the learning process, instead of querying new samples to estimate $L(h,w^{t})$ in each round $t$.
Recall that $\overline{w}^{t}_{i}=\max_{1\leq\tau\leq t}w_{i}^{\tau}$. To ensure the $O(\varepsilon)$-uniform convergence with respect to $w^{t}$ in each round $t$, it suffices to choose $n_{i}^{t}\geq N^{*}\overline{w}_{i}^{t}$ for all $i\in[k]$.
4.2 Measuring the trajectory of Hedge
Our second technical point is to bound the maximal point of the whole trajectory $\{w^{t}\}_{t=1}^{T}$ of the Hedge algorithm. That is, we aim to control
$\|\overline{w}^{T}\|_{1}=\sum_{i=1}^{k}\max_{1\leq t\leq T}w_{i}^{t}.$
To explain the high-level idea, we consider the learning process in Algorithm 2 with full knowledge of the loss vector.
@@ 4.2 Measuring the trajectory of Hedge
2 Initialization: $T=\frac{100\log(k)}{\varepsilon^{2}}$, $\eta=\frac{1}{10}\varepsilon$, $W^{1}_{i}=1$ for all $1\leq i\leq k$.
@@
4       $w_{i}^{t}\leftarrow\frac{W_{i}^{t}}{\sum_{j}W_{j}^{t}}$ for all $1\leq i\leq k$.
5       $y^{t}\leftarrow\arg\min_{y\in\mathcal{Y}}(w^{t})^{\top}y$.
6       Update $W_{i}^{t+1}\leftarrow W_{i}^{t}\cdot\exp(\eta y_{i}^{t})$.

### v3

@@ 2 Problem formulation
This section formulates the multi-distribution learning problem. We assume throughout that each datapoint takes the form of $(x,y)\in\mathcal{X}\times\mathcal{Y}$, with $\mathcal{X}$ (resp. $\mathcal{Y}$) the instance space (resp. label space).
@@
The problem setting encompasses several elements below.
  • Hypothesis class. Suppose we are interested in a hypothesis class $\mathcal{H}$, comprising a set of candidate functions from the instance space $\mathcal{X}$ to the label space $\mathcal{Y}$. Overloading the notation, we use $h_{\pi}$ to represent a randomized hypothesis associated with a probability distribution $\pi\in\Delta(\mathcal{H})$, meaning that a hypothesis $h$ from $\mathcal{H}$ is randomly selected according to distribution $\pi$. When it comes to a VC class, we assume that the hypotheses in $\mathcal{H}$ are binary-valued, and that the VC dimension (vapnik1994measuring) of $\mathcal{H}$ is
    $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d.$ (6)
@@ 3 Algorithm
In this section, we present our proposed algorithm for learning VC classes. Before proceeding, we find it convenient to introduce some notation concerning the loss under mixed distributions. Specifically, for any distribution $w=[w_{i}]_{1\leq i\leq k}\in\Delta(k)$ and any hypothesis $h\in\mathcal{H}$, the risk over the mixture $\sum_{i\in[k]}w_{i}\mathcal{D}_{i}$ of data distributions is denoted by:
@@
2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}\coloneqq\frac{4000\left(k\log(k/\varepsilon_{1})+d\log(kd/\varepsilon_{1})+\log(1/\delta)\right)}{\varepsilon_{1}^{2}}.$
@@
       /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}$ only if $w^{t}$ changes sufficiently. */
@@ 9                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
12       compute $h^{t}\leftarrow\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})$, where
$\widehat{L}^{t}(h,w^{t})\coloneqq\sum_{i=1}^{k}\frac{w^{t}_{i}}{n_{i}^{t}}\cdot\sum_{j=1}^{n_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (9)
@@
15             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
@@
17
output: a randomized hypothesis $h^{\mathsf{final}}$ uniformly distributed over $\{h^{t}\}_{t=1}^{T}$.
Algorithm 1 Hedge for multi-distribution learning on VC classes ($\mathtt{MDL}\text{-}\mathtt{Hedge}\text{-}\mathtt{VC}$)
Remark 1.
Note that in Algorithm 1, the quantities $\{n_{i}^{t}\}_{1\leq t\leq T}$ are designed to be non-decreasing in $t$.
Remark 2.
In line 1 of Algorithm 1, we also allow $h^{t}$ to be $\varepsilon_{1}$-best response that obeys $\widehat{L}^{t}(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})+\varepsilon_{1}$; our theory remains unchanged if we make this modification.
Following the game dynamics proposed in previous works (awasthi2023sample; haghtalab2022demand), our algorithm alternates between computing the most favorable hypothesis (performed by the learner) and estimating the most challenging mixture of data distributions (performed by the adversary), with the aid of no-regret learning algorithms (roughgarden2016twenty; shalev2012online). More specifically, in each round $t$, our algorithm performs the following two steps:
@@
    Given hypothesis $h^{t}$, we compute an updated weight vector $w^{t+1}\in\Delta(k)$ — and hence an updated mixed distribution $\mathcal{D}^{(t+1)}=\sum_{i\in[k]}w_{i}^{t+1}\mathcal{D}_{i}$. The weight updates are carried out using the celebrated Hedge algorithm (freund1997decision) designed for online adversarial learning, [^5] Note that the Hedge algorithm is closely related to Exponentiated Gradient Descent, Multiplicative Weights Update, Online Mirror Descent, etc (arora2012multiplicative; shalev2012online; hazan2022introduction). in an attempt to achieve low regret even when the loss vectors are adversarially chosen. More precisely, we run
@@
In words, the min-player and the max-player in our algorithm follow the best-response dynamics and no-regret dynamics, respectively (note that the Hedge algorithm is known to be a no-regret algorithm (roughgarden2016twenty)). At the end of the algorithm, we output a randomized hypothesis $h^{\sf final}$ that is uniformly distributed over the hypothesis iterates $\{h^{t}\}_{1\leq t\leq T}$ over all $T$ rounds, following common practice in online adversarial learning.
While the above paradigm has been adopted in past works (awasthi2023sample; haghtalab2022demand), the resulting sample complexity depends heavily upon how data samples are collected and utilized throughout the learning process. For instance, awasthi2023sample — which also alternates between best response and no-regret algorithm — draws fresh data at each step of every round, in order to ensure reliable estimation of the loss function of interest through elementary concentration inequalities. This strategy, however, becomes wasteful over time, constituting one of the main sources of its sample sub-optimality.
@@
  • Sample reuse in Step (a). In stark contrast to awasthi2023sample that draws new samples for estimating each $h^{t}$, we propose to reuse all samples collected in Step (a) up to the $t$-th round to assist in computing $h^{t}$. As will be made precise in lines 1-1 of Algorithm 1, we shall maintain a growing dataset $\mathcal{S}$ for conducting Step (a) throughout, ensuring that there are $n_{i}^{t}$ samples drawn from distribution $\mathcal{D}_{i}$ in the $t$-th round. These datapoints are employed to construct an empirical loss estimator $\widehat{L}^{t}(h,w^{t})$ for each $h\in\mathcal{H}$ in each round $t$, with the aim of achieving uniform convergence $|\widehat{L}^{t}(h,w^{t})-L(h,w^{t})|\leq O(\varepsilon)$ over all $h\in\mathcal{H}$. More detailed explanations are provided in Section 4.1.
  • Weighted sampling for Step (b). As shown in line 1 of Algorithm 1, in each round $t$, we sample each $\mathcal{D}_{i}$ a couple of times to compute the empirical estimator for $\mathbb{E}_{(x,y)\in\mathcal{D}_{i}}\big{[}\ell(h^{t},(x,y))\big{]}$, where the number of samples depends upon the running weights $\{w_{i}^{\tau}\}$. More precisely, we collect $\left\lceil k\overline{w}_{i}^{t}\right\rceil$ fresh samples from each $\mathcal{D}_{i}$, where $\overline{w}_{i}^{t}:=\max_{1\leq\tau\leq t}w_{i}^{\tau}$ is the maximum weight assigned to $\mathcal{D}_{i}$ up to now. Informally speaking, this strategy is chosen carefully to ensure reduced variance of the estimators given a sample size budget. The interested reader is referred to Section 4.2 and Lemma 17 for more detailed explanations.
@@
4 A glimpse of key technical novelties
In this section, we highlight two technical novelties that empower our analysis: (i) uniform convergence of the weighted sampling estimator that allows for sample reuse (see Section 4.1), and (ii) tight control of certain $\|\cdot\|_{1,\infty}$ norm of the iterates $\{w^{t}\}_{1\leq t\leq T}$ that dictates the sample efficiency (see Section 4.2).
4.1 Towards sample reuse: uniform concentration and a key quantity
Recall that in Algorithm 1, we invoke the empirical risk estimator $\widehat{L}^{t}(h,w^{t})$ as an estimate of the true risk of hypothesis $h$ over the weighted distribution specified by $w^{t}$ (cf. (9)). In order to facilitate sample reuse when constructing such risk estimators across all iterations, it is desirable to establish uniform concentration results to control the errors of such risk estimators throughout the execution of the algorithm. Towards this end, our analysis strategy proceeds as follows.
Step 1: concentration for any fixed set of parameters.
Consider any given set of integers $n=\{n_{i}\}_{i=1}^{k}$ and any given vector $w\in\Delta(k)$. Suppose, for each $i\in[k]$, we have $n_{i}$ i.i.d. samples drawn from $\mathcal{D}_{i}$ — denoted by $\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}$ — and let us look at the empirical risk estimator,
$\widehat{L}_{n}(h,w)\coloneqq\sum\nolimits_{i=1}^{k}w_{i}\cdot\frac{1}{n_{i}}\sum\nolimits_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)},$ (12)
which is a sum of independent random variables. Evidently, for a given hypothesis $h$, the variance of $\widehat{L}_{n}(h,w)$ is upper bounded by
$\mathsf{Var}\big{(}\widehat{L}_{n}(h,w)\big{)}\leq\sum\nolimits_{i=1}^{k}\frac{w_{i}^{2}}{n_{i}}\leq\left(\sum\nolimits_{i=1}^{k}w_{i}\right)\frac{1}{\min_{i}n_{i}/w_{i}}=\frac{1}{\min_{i}n_{i}/w_{i}}.$
Assuming that the central limit theorem is applicable, one can derive
@@ 4.1 Towards sample reuse: uniform concentration and a key quantity
$\mathbb{P}\left\{\big{|}\widehat{L}_{n}(h,w)-L(h,w)\big{|}\geq\varepsilon\right\}\lesssim\exp\bigg{(}-\frac{\varepsilon^{2}}{2\mathsf{Var}\big{(}\widehat{L}_{n}(h,w)\big{)}}\bigg{)}\lesssim\exp\bigg{(}-\frac{\varepsilon^{2}}{2}\min_{i}\frac{n_{i}}{w_{i}}\bigg{)}.$
Armed with this result, we can extend it to accommodate all $h\in\mathcal{H}$ through the union bound. For a VC class with $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d$, the celebrated Sauer–Shelah lemma (wainwright2019high, Proposition 4.18) tells us that the set of hypotheses can be effectively compressed into a subset with cardinality no larger than $\exp\big{(}\widetilde{O}(d)\big{)}$. Taking the union bound then yields
@@
$\mathbb{P}\left(\max_{h\in\mathcal{H}}\big{|}\widehat{L}_{n}(h,w)-L(h,w)\big{|}\geq\varepsilon\right)\lesssim\exp\left(\widetilde{O}(d)-\frac{\varepsilon^{2}}{2}\min_{i}\frac{n_{i}}{w_{i}}\right).$
Step 2: uniform concentration.
Next, we would like to extend the above result to establish uniform concentration over all $n$ and $w$ of interest. Towards this, we shall invoke the union bound as well as the standard epsilon-net arguments. Let the set $\mathcal{X}\subseteq\Delta(k)$ be a proper discretization of $\Delta(k)$, with cardinality $\exp\big{(}\widetilde{O}(k)\big{)}$. In addition, given the trivial upper bound $n_{i}\leq T_{1}$ for all $i\in[k]$, we know that there exist at most $T_{1}^{k}=\exp\big{(}\widetilde{O}(k)\big{)}$ possible combinations of $\{n_{i}\}_{i\in[k]}$. We can then apply the union bound to show that
$\mathbb{P}\left\{\exists w\in\mathcal{X}\text{ and feasible }n\text{ s.t.}\big{|}\widehat{L}_{n}(h,w)-L(h,w)\big{|}\geq\varepsilon\right\}$ $\lesssim\exp\bigg{(}\widetilde{O}(k)+\widetilde{O}(d)-\frac{\varepsilon^{2}}{2}\min_{i}\frac{n_{i}}{w_{i}}\bigg{)}.$ (13)
When the discretized set $\mathcal{X}$ is chosen to have sufficient resolution, we can straightforwardly employ the standard covering argument to extend the above inequality to accommodate all $w\in\Delta(k)$ of interest.
Key takeaways.
@@
The above arguments reveal the following high-probability property: whenever we collect $n=\{n_{i}\}_{i=1}^{k}$ samples in the learning process, we could obtain $\varepsilon$-approximation $\widehat{L}_{n}(h,w)$ (see (12)) of $L(h,w)$ for all $h\in\mathcal{H}$ and all $w\in\Delta(k)$ with high probability, provided that
$\min_{i}\frac{n_{i}}{w_{i}}\gtrsim\widetilde{O}\bigg{(}\frac{k+d}{\varepsilon^{2}}\bigg{)}.$ (14)
This makes apparent the pivotal role of the quantity $\min_{i}n_{i}/w_{i}$. In our algorithm, we design the update rule (cf. line 1 of Algorithm 1) to guaranteed that
$\min_{i}\frac{n_{i}^{t}}{w_{i}^{t}}\gtrsim T_{1}\geq\widetilde{\Omega}\bigg{(}\frac{k+d}{\varepsilon^{2}}\bigg{)}$ (15)
for all $1\leq t\leq T$. In fact, this explains our choice of $T_{1}$ in Algorithm 1. Crucially, the aforementioned uniform concentration result allows us to reuse samples throughout the learning process instead of drawing fresh samples to estimate $L(h,w^{t})$ in each round $t$ (note that the latter approach clearly falls short of data efficiency). To conclude, to guarantee $\varepsilon$-uniform convergence for all rounds, it suffices to choose $T_{1}=\widetilde{\Omega}\big{(}\frac{k+d}{\varepsilon^{2}}\big{)}$.
Finally, recall that $n_{i}^{t}\asymp T_{1}\overline{w}_{i}^{t}$ for each $i\in[k]$ and $t\leq T$, with $\overline{w}^{t}_{i}\coloneqq\max_{1\leq\tau\leq t}w_{i}^{\tau}$; taking $n_{i}^{t}\asymp T_{1}\overline{w}_{i}^{t}$ (as opposed to $n_{i}^{t}\asymp T_{1}w_{i}^{t}$) ensures that the sample size $n_{i}^{t}$ is monotonically non-decreasing in $t$. With (15) in mind, the total number of samples collected within $T$ rounds in Algorithm 1 obeys
$\frac{1}{T_{1}}\sum\nolimits_{i=1}^{k}n_{i}^{T}\asymp\sum\nolimits_{i=1}^{k}\overline{w}_{i}^{T}\eqqcolon\|\overline{w}^{T}\|_{1}.$ (16)
This threshold $\|\overline{w}^{T}\|_{1}$ — or equivalently, the $\|\cdot\|_{1,\infty}$ norm of $\{w^{t}\}_{1\leq t\leq T}$ — is a critical quantity that we wish to control. In particular, in the desirable scenario where $\|\overline{w}^{T}\|_{1}\leq\widetilde{O}(1)$, the total sample size obeys $\sum\nolimits_{i=1}^{k}n_{i}^{T}\asymp T_{1}\|\overline{w}^{T}\|_{1}=\widetilde{O}\big{(}\frac{k+d}{\varepsilon^{2}}\big{)}$.
4.2 Bounding the key quantity $\|\overline{w}^{T}\|_{1}$ by tracking the Hedge trajectory
Perhaps the most innovative (and most challenging) part of our analysis lies in controlling the $\|\cdot\|_{1,\infty}$ norm of $\{w_{i}^{t}\}_{1\leq t\leq T}$, whose critical importance has been pointed out in Section 4.1.
Towards this end, the key lies in carefully tracking the dynamics of the Hedge algorithm. To elucidate the high-level idea, let us look at a simpler minimax optimization problem w.r.t. the set of loss vectors in the convex hull of a set $\mathcal{Y}\subseteq\mathbb{R}^{k}$:
where the equivalence arises from von Neumann’s minimax theorem (v1928theorie). Consider the following algorithm (cf. Algorithm 2) tailored to this minimax problem, assuming perfect knowledge about the loss vectors. [^6] Note that in Algorithm 1, we can only estimate the loss vector using the collected samples. Additional efforts are needed to reduce the variability (see line 1 in Algorithm 1).
@@
2 Initialization: $T=\frac{100\log(k)}{\varepsilon^{2}}$, $\eta=\frac{1}{10}\varepsilon$, and $W^{1}_{i}=1$ for all $1\leq i\leq k$.
@@
4       compute $w_{i}^{t}\leftarrow\frac{W_{i}^{t}}{\sum_{j}W_{j}^{t}}$ for every $1\leq i\leq k$.
5       compute $y^{t}\leftarrow\arg\min_{y\in\mathcal{Y}}~{}\langle w^{t},y\rangle$.
6       update $W_{i}^{t+1}\leftarrow W_{i}^{t}\exp(\eta y_{i}^{t})$ for every $1\leq i\leq k$.
7